Hi, I need to solve this wood truss using method of joints to find every force for each member and whether it is in tension or compression. I also need to determine the reaction forces at both pinned connections.SY - TY= 1940 * 3,740 * 6Y - 5Y - 3,740 * 3,740 44 = 3Y = # 3,740 2Y = 3,740 * IY = 3,740 # 1,940 * > 4-0 1.4-0 1.4-0.
Answer #1

1990 3140 3140 1940. 3740 3740 3740 3740 u D H I 4 1875 VM B F g A n RB В RA tan o = BM BN. RA + Rp = 26.320 BM = Bn tom (8.5ty pont m 3740 70.548 FML к B S 2o > FAN a 13 116.2 FBM = 13160 FEM. Fore Sino FILM Sing 13116.2 Smo - 3740 +13160 0.1478 FomE E 8 jont k FILE 3740 18750.65 1 mich LI=1.1956. 0= 16.000 Fran K. PREB 212 97.9 FIL Sin 16.64 + 3140 + 18450.65 = FER =-785FuL = 282711. Tension) 9740 jont H FUE → FHLI =75703 6.97 Full THE = 157036.97 (Tension) FHG = 3740 (compression) E 1:999 1.5FAE =745390.6+ FoE Cos (85) Cos (36.77) (FOE For Sin (8.5) + FoE Cos (8.57-645390.6 6166.77) sin (36.77) = 38420.8. Sin (9.5)